OpenStudy (jacksonjrb):

Help me please. Algebra 2/Trig

OpenStudy (jacksonjrb):

Find the solutions for the equation. \[3x^3-x^2-9x+3=0\]

OpenStudy (jacksonjrb):

I've used Descartes Rule and the Rational Root Theorem already.

OpenStudy (jacksonjrb):

I know that there are 2 or 0 possible positive roots and 1 possible negative roots and that the possible roots are +/- 1, +/-3, and +/-\[\frac{ 1 }{ 3 }\]

OpenStudy (anonymous):

3x^3-x^2-9x+3 = 0 (3x-1)(x^2 - 3) = 0 So 3x-1 = 0 x = 1/3 or x^2 - 3 = 0 x^2 = 3 x = sqrt(3)

OpenStudy (jacksonjrb):

I tried using synthetic division but the problem seems to be a bit weird.

OpenStudy (jacksonjrb):

So the solutions are 1/3 and sqrt(3)?

OpenStudy (jacksonjrb):

There is one negative roots though...

OpenStudy (anonymous):

yes

OpenStudy (jacksonjrb):

*root

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sqrt (3) is 1.73 or -1.73

OpenStudy (anonymous):

the solusions are 1/3, 1.73, and -1.73

OpenStudy (jacksonjrb):

Ah, that makes sense.

OpenStudy (lyrae):

\[x^2 - 3 = 0 \rightarrow x = \pm \sqrt{3}\]

OpenStudy (jacksonjrb):

Thank you.

OpenStudy (anonymous):

or the solusions are the solusions are 1/3, sqrt(3), and -sqrt(3)

OpenStudy (anonymous):

you are welcome :)